OpenStudy (anonymous):

Evaluate 3x - 2 if x = 5

OpenStudy (anonymous):

f(5)=3(5)-2 f(5)=15-2 f(5)=13

OpenStudy (anonymous):

3(5) -2 = 15 -2 = 13

OpenStudy (anonymous):

15